RC Meeting, Thursday, January 20, 2022
RC Meeting Information
RC Daily Meetings are at 10 AM, remote
- Zoom remote connection information:
Join ZoomGov Meeting https://jlab-org.zoomgov.com/j/1607100885?pwd=ZFBsRTFsYm5Zb0laZ0tyZXNjcEE0QT09 Meeting ID: 160 710 0885 Passcode: 595071 One tap mobile +16692545252,,1607100885# US (San Jose) +16468287666,,1607100885# US (New York) Dial by your location +1 669 254 5252 US (San Jose) +1 646 828 7666 US (New York) +1 669 216 1590 US (San Jose) +1 551 285 1373 US 833 568 8864 US Toll-free Meeting ID: 160 710 0885 Find your local number: https://jlab-org.zoomgov.com/u/ab2CekLXgQ
- Fall 2021 Shift Schedule
- PionLT Physics Status
- PionLT Runplans
- Hall-C Controls
- Hall-C Logbook
- MCC White Board
- OPS 7:45 RC meeting Zoom link
- OPS daily 8:00 meeting and Wed 1:30 new Zoom link
- Zoom link for remote shifts: https://jlab-org.zoomgov.com/j/1613876340?pwd=WG1pOFVsaXdGaHJvbG15QVBjL3k5Zz09
- Social distancing requirements back in place. Please maintain a separation of 6 feet during shifts.
- See updated guidance on JLab COVID-19 isolation and quarantine protocols: https://www.jlab.org/memo/jlab-covid-19-isolation-and-quarantine-requirements
- Communicate with Occupational Medicine if you have COVID-19 symptoms, have tested positive for COVID-19, or have been exposed to someone who has tested positive for COVID-19: (email@example.com or firstname.lastname@example.org).
- Please keep an eye on the HMS NMR during shift. Due to the NMR locking instability, the alarm has been been disabled, so slow drifts of the NMR signal may be missed.
- See note on lower right TV screen for bringing up Zoom in CH for remote 3rd shifts
Current RC: Simona Malace
- Q2=3.85, W=3.07, low epsilon, SHMS left setting, finished running at this Q2. The pion count on LD2 was 10038 right before the controlled access.
- At the beginning of the shift start the switch to Q2 = 5, W2 = 2.95, low epsilon
- Start Q2 = 5
- During day shift make an access to set SHMS angle to 6.17 deg (Hall C techs will watch rotation from down in the Hall while Steve L. will watch rotation remotely - cannot be on site due to possible bad weather).
- Run at Q2 = 5, SHMS angle = 6.17 deg on LH2 and then Al
- Run at Q2 = 5, SHMS angle = 8,17 deg on LH2 and Al (SHMS rotation from 6.17 deg to 8.17 deg will be done remotely but will be watched - also remotely - by Steve L. or Amy - TBD)
Current Run Plan
- Start Q2=5, W=2.95, low epsilon on LH2
- SHMS settings: theta = 6.17 degrees, P=6.719 GeV
- HMS settings: theta = 44.37 degree, P=-1.104 GeV
- Targets = LH2, aluminum dummy
Notices in Effect for Shift Crew
- The Raster Coil pair Ax and Ay are off.
Raster needs to be at 4x4.Raster gui has been updated - can now request 2x2.
- With LHe alarms it is good to check the cameras for plumes. If no plume and it is below the LO alarm and above LOLO then monitor with a plot and halog the problem.
- With LNitrogen alarms are not that critical. Just need to start monitoring the problem with plot and Halog.
- HMS NMR is not locking stably - the alarm is turned off. Please check on the NMR readback (while it is locked) intermittently to make sure the HMS dipole field is not drifting,
- Vertical beam position on target has changed a little: Dave Meekins is not concerned about an offset of this size (<0.5 mm).
- There is a rotation issue that has come up if moving HMS between 19 to 25 degrees. It usually gets stuck around this range, please be wary of this issue. We will likely skip settings in this range if this drive controller error cannot be solved. For singles it will not be a big deal to skip settings for HMS at those angles. We need a longer access to address this issue.
- The Raster Coil pair Ax and Ay are off
- Diffuser limit at 46uA (Note: is this still valid?)
Crew Chiefs, We have permission to raise the Hall C diffuser raster limit to allow running >46uA to Hall C. This limit is normally imposed when one of the two rasters is not operative. Since the hall was able to double the running raster current to get the same size raster on the target as having both rasters, Keith is allowing us to raise the limit. To get to the setting, open the Hall C diffuser interlocks screen. Go to the bottom right corner and open the Expert screen. Just below and to the right of the FSD bits, raise the "Raster 1 OK Limit" to slightly above the highest current the hall wants to run at. Do not change the Raster 2 limit. Make a log entry with screen shots. Make an ABIL entry, System MPS and an expected removal date of 3 Jan. 2022. Paul
Planned Accesses to the Hall
- Radcon will arrive around 7:30 am to start the full survey
- Hall C techs and Heidi will access the hall hopefully around 8:30 am for: SHMS rotation to 6.17 deg and to fix SHMS Q1 lead-flow controller
Planned Accelerator Configuration Changes
- Hall B pass change planned for Monday, Jan. 24
- 2L18 cryocycle required for next run period - planned for next Tuesday.